# Multistep Consistency Models

### v2

@@ 5.1 Quantitative Evaluation on ImageNet
iCT-deep (Song & Dhariwal, 2023) 1 3.25
@@
For our ImageNet experiments we trained diffusion models on ImageNet64 and ImageNet128 in a base and large variant. We initialize the consistency models from the pre-trained diffusion model weights which we found to greatly increase robustness and convergence. Both consistency training and distillation are used. Classifier Free Guidance (Ho & Salimans, 2022) was used only on the base ImageNet128 experiments. For all other experiments we did not use guidance because it did not significantly improve the FID scores of the diffusion model. All consistency models are trained for $200,000$ steps with a batch size of $2048$ and an step schedule that anneals from $64$ to $1280$ in $100.000$ steps with an exponential schedule.
In Table 1 it can be seen that as the multistep count increases from a single consistency step, the performance considerably improves. For instance, on the ImageNet64 Base model it improves from 7.2 for one-step, to 2.7 and further to 1.8 for two and four-steps respectively. In there are generally two patterns we observe: As the steps increase, performance improves. This validates our hypothesis that more steps give a helpful trade-off between sample quality and speed. It is very pleasant that this happens very early: even on a complicated dataset such as ImageNet128, our base model variant is able achieve 2.1 FID in 8 steps, when consistency distilling.
To draw a direct comparison between Progressive Distillation (PD) (Salimans & Ho, 2022) and our approaches, we reimplement PD using aDDIM and we use same base architecture, as reported in Table 3. With our improvements, PD can attain better performance than previously reported in literature. However, compared to MultiStep CT and CD it starts to degrade in sample quality at low step counts. For instance, a 4-step PD model attains an FID of 2.4 whereas CD achieves 1.7.
Further we ablate whether annealing the step schedule is important to attain good performance. As can be seen in Table 2, it is especially important for low multistep models to anneal the schedule. In these experiments, annealing always achieves better performance than tests with constant steps at $128,256,1024$. As more multisteps are taken, the importance of the annealing schedule decreases.
@@
Compared to existing works in literature, we achieve SOTA FID scores in both ImageNet64 on 2-step, 4-step and 8-step generation. Interestingly, we achieve approximately the same performance using single step CD compared to iCT-deep (Song & Dhariwal, 2023), which achieves this result using direct consistency training. Since direct training has been empirically shown to be a more difficult task, one could conclude that some of our hyperparameter choices may still be suboptimal in the extreme low-step regime. Conversely, this may also mean that multistep consistency is less sensitive to hyperparameter choices.
In addition, we compare on ImageNet128 to our reimplementation of Progressive Distillation. Unfortunately, ImageNet128 has not been widely adopted as a few-step benchmark, possibly because a working deterministic sampler has been missing until this point. For reference we also provide the recent result from (Kingma & Gao, 2023). Further, with these results we hope to put ImageNet128 on the map for few-step diffusion model evaluation.
5.2 Qualitative Evaluation on Text to Image modelling
In addition to the analysis on ImageNet, we study the effects on a text-to-image models. We distill a 16-step consistency model a base teacher model. In Table 5 one can see that Multistep CD is able to distill its teacher almost perfectly in terms of FID. The loss of clip score can be attributed to the guidance distillation, which a baseline 256-step student model also has trouble distilling. Compared to the guidance-distilled baseline, the 16-CD model has no loss in performance measured in CLIP and FID on the low guidance setting (and for the high guidance setting only a minor degradation in FID). Even the 8-step CD model attains an impressive FID score of 8.1, which is well below the existing literature.
In Figure 2 and 5 we compare samples from our 16-step CD aDDIM distilled model to the original 100-step DDIM sampler. Because the random seed is shared we can easily compare the samples between these models, and we can see that there are generally minor differences. In our own experience, we often find certain details more precise, at a slight cost of overall construction. Another comparison in Figure 4 shows the difference between a DDIM distilled model (equivalent to $\eta=0$ in aDDIM) and the standard DDIM sampler. Again we see many similarities when sharing the same initial random seed.
@@ 6 Conclusions
Broader Impacts
This paper proposes a method to speed up sampling from diffusion models. Although generative models may be used for positive applications such as enhancing human creativity or drug discovery, they may also be used to create deepfakes or misinformation. Hence, enabling faster sampling may amplify both the positive and the negative impacts of generative modelling.
@@ A.1 Setup
In this paper we follow the setup from simple diffusion (Hoogeboom et al., 2023). Following their approach, we use a standard UViTs. These are UNets with MLP blocks instead of convolutional layers when a block has self-attention, making the entire block a transformer block. This contains the details for the architecture and how to define diffusion process. There are some minor specifics which we share per experiment below.
Multistep Constinency Hyperparameters
For all ImageNet runs (small/large, 1 through 16 step) we use a log-linear interpolated schedule from 64 teacher steps to 1280 teacher steps, annealed over 100000 training iterations at a batch size of 2048. We use a xvar_frac of 0.75 for aDDIM. And we use a huber epsilon of 1e-4. The model is trained for 200000 steps. The interpolation starts quite low and takes a long time, and these settings are somewhat excessive for the larger student step models such as the 8- or 16-step model. However, fixing these settings for the model allowed for clean comparisons.
@@
For the ImageNet64 experiments, the levels of the UViT small are as follows. Down: 3 ResNet blocks with 256 channels, 3 Transformer Blocks with 512 channels both stages ending with an average pool. Middle: 16 transformer blocks with 1024 channels, mlp expansion factor is 4. Up, matching the down blocks, starting a stage with a nearest neighbour upsampling and obviously no pooling. Dropout is applied to the middle with a factor of 0.2. For the large variant, all channels are multiplied by 2, and dropout is applied to all transformers albeit with a lower factor of 0.1. The network is trained with an interpolated cosine schedule from noise resolution 32 to 64 at a resolution of 64 (this is practically identical to a normal cosine schedule).
@@
For the ImageNet128 experiments, the UViT is the same as the UViT for ImageNet64, but with an extra 3 ResNet Blocks at the resolution 128x128 with 128 channels at both the start and the end of the UViT. For completeness, down: 3 ResNet blocks with 128 channels, 3 ResNet blocks with 256 channels, 3 Transformer Blocks with 512 channels both stages ending with an average pool. Middle: 16 transformer blocks with 1024 channels, mlp expansion factor is 4. Up, matching the down blocks, starting a stage with a nearest neighbour upsampling and no pooling.
@@ A.3 Datasets
A.4 Additional results

### v3

@@ 5.1 Evaluation on ImageNet
iCT-deep (Song and Dhariwal, 2023) 1 3.25
@@
To draw a direct comparison between Progressive Distillation (PD) (Salimans and Ho, 2022) and our approaches, we reimplement PD using aDDIM and we use same base architecture, as reported in Table 3. With our improvements, PD can attain better performance than previously reported in literature. However, compared to MultiStep CT and CD it starts to degrade in sample quality at low step counts. For instance, a 4-step PD model attains an FID of 2.4 whereas CD achieves 1.7.
In Tbl. 4 we ablate the effect of using adjusted DDIM as a teacher. Empirically, we observe that the adjusted sampler is important when more student steps are used. In contrast, vanilla DDIM works better when few steps are taken and the student does not get close to the teacher as measured in FID.
Further we ablate whether annealing the step schedule is important to attain good performance. As can be seen in Tbl. 2, it is especially important for low multistep models to anneal the schedule. In these experiments, annealing always achieves better performance than tests with constant teacher steps at $128,256,1024$. As more student steps are taken, the importance of the annealing schedule decreases.
@@
Compared to existing works in literature, we achieve SOTA FID scores in both ImageNet64 and Imagenet128 with 4-step and 8-step generation. Interestingly, we achieve approximately the same performance using single step CD compared to iCT-deep (Song and Dhariwal, 2023), which achieves this result using direct consistency training. Since direct training has been empirically shown to be a more difficult task, one could conclude that some of our hyperparameter choices may still be suboptimal in the extreme low-step regime. Conversely, this may also mean that multistep consistency is less sensitive to hyperparameter choices.
In addition, we compare on ImageNet128 to our reimplementation of Progressive Distillation. Unfortunately, ImageNet128 has not been widely adopted as a few-step benchmark, possibly because a working deterministic sampler has been missing until this point. For reference we also provide the recent result from (Kingma and Gao, 2023). Further, with these results we hope to put ImageNet128 on the map for few-step diffusion model evaluation.
5.2 Evaluation on Text to Image modelling
In addition to the analysis on ImageNet, we study the effects on text-to-image models. We distill a 16-step consistency model from a base teacher model. In Table 6 one can see that Multistep CD is able to distill its teacher almost perfectly in terms of FID. The loss of clip score can be attributed to the guidance distillation, which a baseline 256-step student model also has trouble distilling. Compared to the guidance-distilled baseline, the 16-CD model has no loss in performance measured in CLIP and FID on the low guidance setting (and for the high guidance setting only a minor degradation in FID). Even the 8-step CD model attains an impressive FID score of 8.1, which is well below the existing literature.
In Figure 2 and 6 we compare samples from our 16-step CD aDDIM distilled model to the original 100-step DDIM sampler. Because the random seed is shared we can easily compare the samples between these models, and we can see that there are generally minor differences. In our own experience, we often find certain details more precise, at a slight cost of overall construction. Another comparison in Figure 4 shows the difference between a DDIM distilled model (equivalent to $\eta=0$ in aDDIM) and the standard DDIM sampler. Again we see many similarities when sharing the same initial random seed.
@@ A.1 Setup
In this paper we follow the setup from simple diffusion (Hoogeboom et al., 2023). Following their approach, we use a standard UViTs. These are UNets with MLP blocks instead of convolutional layers when a block has self-attention, making the entire block a transformer block. This contains the details for the architecture and how to define diffusion process. There are some minor specifics which we share per experiment below. All runs are initialized using the parameters of a pretrained diffusion models.
Multistep Consistency Hyperparameters
For all ImageNet runs (small/large, 1 through 16 step) we use a log-linear interpolated schedule from 64 teacher steps to 1280 teacher steps, annealed over 100000 training iterations which means:
$\small{N_{\mathrm{teacher}}(i)=\exp(\log 64+\operatorname{clip}(i/100.000,0,1)\cdot(\log 1280-\log 64))}$
. The batch size is 2048. We use a xvar_frac of 0.75 for aDDIM. And we use a huber epsilon of 1e-4. The model is trained for 200000 steps. The interpolation starts quite low and takes a long time, and these settings are somewhat excessive for the larger student step models such as the 8- or 16-step model. However, fixing these settings for the model allowed for clean comparisons. These runs anneal the teacher steps using
@@
For the ImageNet64 experiments, the levels of the UViT small are as follows. Down: 3 ResNet blocks with 256 channels, 3 Transformer Blocks with 512 channels both stages ending with an average pool. Middle: 16 transformer blocks with 1024 channels, mlp expansion factor is 4. Up, matching the down blocks, starting a stage with a nearest neighbour upsampling and obviously no pooling. Dropout is applied to the middle with a factor of 0.2. For the large variant, all channels are multiplied by 2, and dropout is applied to all transformers albeit with a lower factor of 0.1. The network is trained with an interpolated cosine schedule from noise resolution 32 to 64 at a resolution of 64 (this is practically identical to a normal cosine schedule). The small and large model have $394$M and $1.23$B parameters, respectively.
@@
For the ImageNet128 experiments, the UViT is the same as the UViT for ImageNet64, but with an extra 3 ResNet Blocks at the resolution 128x128 with 128 channels at both the start and the end of the UViT. For completeness, down: 3 ResNet blocks with 128 channels, 3 ResNet blocks with 256 channels, 3 Transformer Blocks with 512 channels both stages ending with an average pool. Middle: 16 transformer blocks with 1024 channels, mlp expansion factor is 4. Up, matching the down blocks, starting a stage with a nearest neighbour upsampling and no pooling. The small and large model have $397$M and $1.25$B parameters, respectively.
@@ A.3 Datasets
A.4 Teacher sampling
Figure 5: Comparison of different sampling methods for the cosine schedule (left) and the sigma schedule used by Karras et al. (2022) (right) on Imagenet64. Note that aDDIM with a (shifted) cosine schedule is the best performing model overall except for the 64 function evaluation.
Fig. 5 compares various samplers including the 2nd order Heun sampler. Additionally, a stochastic version of DDIM is included (noise DDIM) where we add random Guassian noise directly to the model prediction. This direct noise injection breaks the determinism of DDIM and is therefore not a useful sampler for consistent distillation. However, it behaves very similarly to the aDDIM which seems to indicate that our heuristic noise correction is accurately simulating the positive effects of noise injection in the sampler.
Interestingly, we observe a significant difference in the relative quality of various sampling methods depending on the noise schedule used at evaluation. The Heun sampler favors the schedule introduced by Karras (Karras et al., 2022) while the noisy methods seem to work better with a standard cosine schedule. One possible explanation is that the asymptotic behavior of the cosine schedule favours the noise injection methods. Previous work has indicated that the asymptotic behavior a noise schedule is important to fully capture the data distribution (Lin et al., 2024). We consider investigating the interaction between schedules and samplers and interesting opportunity for future work.
A.5 Additional results